# Combined Fit of Spectrum and Composition for FR0 Radio Galaxy Emitted Ultra-High-Energy Cosmic Rays with Resulting Secondary Photons and Neutrinos

### v1

@@ Abstract
This study comprehensively investigates the gamma-ray dim population of Fanaroff-Riley Type 0 (FR0) radio galaxies as potentially significant sources of ultra-high-energy cosmic rays (UHECRs, E 1018 eV) detected on Earth. While individual FR0 luminosities are relatively low compared to the more powerful Fanaroff-Riley Type 1 and Type 2 galaxies, FR0s are substantially more prevalent in the local universe, outnumbering the more energetic galaxies by a factor of 5 within a redshift of z 0.05.
Employing CRPropa3 simulations, we estimate the mass composition and energy spectra of UHECRs originating from FR0 galaxies for energies above 1018.6 eV. This estimation fits data from the Pierre Auger Observatory (Auger) using three extensive air shower models; both constant and energy-dependent observed elemental fractions are considered. The simulation integrates an isotropic distribution of FR0 galaxies, extrapolated from observed characteristics, with UHECR propagation in the intergalactic medium, incorporating various plausible configurations of extragalactic magnetic fields, both random and structured. We then compare the resulting emission spectral indices, rigidity cutoffs, and elemental fractions with recent Auger results. In total, 25 combined energy spectrum and mass composition fits are considered.
Beyond the cosmic ray fluxes emitted by FR0 galaxies, this study predicts the secondary photon and neutrino fluxes from UHECR interactions with intergalactic cosmic photon backgrounds. The multi-messenger approach, encompassing observational data and theoretical models, helps elucidate the contribution of low luminosity FR0 radio galaxies to the total cosmic ray energy density.
@@
Ultra-high-energy-cosmic radiation (1733), Fanaroff-Riley radio galaxies (526), Low-luminosity AGN (2033), Extragalactic magnetic fields (507), Cosmic background radiation (317), Gamma-rays (637), Neutrino astronomy (1100), Cosmic ray showers (327)1 Introduction
@@
Ultra-high-energy cosmic rays (UHECRs) are the most energetic particles observed in the Universe, with energies exceeding eV. Their origins are an open astrophysical question, with potential sources including active galactic nuclei (AGN), various radio galaxies, gamma-ray bursts, and starburst galaxies. The Pierre Auger Observatory (Auger), located in Argentina, was established as the largest (3000 km2) UHECR hybrid detector in the world to study these particles [1]. Auger provides the highest statistics data on the energy spectra, composition, and arrival directions of UHECR, offering essential information to explore their sources and propagation mechanisms.
@@
Low luminosity Fanaroff-Riley (FR0 Class) radio galaxies [2] (radio-loud jetted AGN) have emerged as a promising candidate class of sources of UHECRs detected on Earth. Recent surveys have highlighted that FR0s are approximately five times more prevalent in the local universe compared to their more energetic counterparts, Fanaroff-Riley Type 1 (FR1) and Type 2 sources [3], qualifying them as potentially substantial contributors to the UHECR energy density. Despite their similar core radio properties to FR1s [4], FR0s display much weaker extended radio emission; however, their estimated jet power, which significantly exceeds the required cosmic ray power per source [5], suggests they could substantially contribute to the UHECR flux measured on Earth. Studies utilizing the average FR0 spectral energy distribution suggest that these galaxies can accelerate UHECRs to the highest observed energies through hybrid acceleration, combining Fermi-I pre-acceleration with gradual shear acceleration [6]. Therefore, they fulfill both the Hillas criterion [7] and the energetics criteria [8, 9].
@@
Multi-messenger astrophysics, which involves the combined study of various cosmic messengers such as photons, neutrinos, and cosmic rays, offers an even more comprehensive understanding of astrophysical phenomena. This approach helps identify and characterize sources of UHECRs and refine our knowledge of the mechanisms that generate and propagate these particles. Regarding the multi-messenger properties of FR0s, lepto-hadronic jet-disc modeling of core-emission from quiet low luminosity AGN jets predicts that FR0s are expected to be relatively weak gamma-ray and neutrino emitters [10, 11]. This theoretical evaluation of weak multi-messenger emission aligns with the results of searches for gamma-ray emitting FR0 galaxies. For example, stacking analyses were necessary to detect a very small unresolved FR0 population in Fermi-LAT data and identify them as gamma-ray emitters [12, 13]. A recent study [14] using various Fermi-LAT catalogs as sources of cosmic rays in CRPropa simulations concluded that previously resolved (therefore overwhelmingly non-FR0) gamma-ray sources cannot account for the large-scale anisotropy of UHECR events above 8 EeV as established by the Pierre Auger Collaboration [15]. This discrepancy led the authors of [14] to conclude that another population of UHECR sources, either lacking gamma-ray emission or unresolved by current-generation telescopes, must exist.
@@
The effect on expected arrival distribution of diffuse UHECRs in the intergalactic magnetic field has been examined in several studies, e.g., in [16, 17, 18] that provide constraints on the source emission spectra and composition (e.g., [19, 20]). The Auger large angular scale arrival direction analysis points to the extragalactic origins of UHECR for energies E 8 EeV [15]. A recent three-component combined fit of the UHECR energy spectrum, extensive air shower (EAS) maxima distributions (X), and arrival directions supports an astrophysical model primarily comprised of homogeneous background sources and an adaptable addition from nearby source candidates [21]. For instance, if these additional nearby sources are a selection of 44 starburst galaxies, they could contribute up to a 20% flux fraction at 40 EeV [21] — a sizable fraction of the remaining bulk of UHECRs may be from FR0s.
@@
In this study, we leverage the strengths of multi-messenger observations to investigate the properties of FR0 radio galaxies as potential UHECR sources. We simulate isotropic FR0 UHECR emission, extrapolating from measured FR0 properties and propagating them to Earth through plausible intergalactic magnetic fields using CRPropa3 [26]. By a combined fit of these simulations to energy and composition using Pierre Auger Observatory data [27, 28], we estimate the UHECR mass composition and energy spectra emitted by FR0s. These combined energy spectrum and composition fits are done using two nuclei emission composition models - a constant nuclei composition emission over the energy range and an evolving emission that depends on energy. Additionally, we present estimates of the cosmogenic secondary photon and neutrino fluxes resulting from UHECR interactions with cosmic photon backgrounds.
@@ 2 Pierre Auger Observatory Data
Auger performs high-precision and large-statistics analyses of energy spectra, mass composition, and arrival directions of UHECR data which are essential for assessing theoretical models of cosmic ray origins and propagation. This study utilizes the energy spectrum and composition datasets to systematically extrapolate the observed characteristics of UHECRs to predictions of FR0 galaxy emission. The Auger data used in fitting simulations of FR0 radio galaxy UHECR emission is the publicly available energy spectrum from [28] and composition from [27] (using the method of [29]) for energies E 1018.6 eV. These results using 13 years of data are shown in Figure 6. The composition is given in terms of mean log mass number , which is evaluated from fluorescence detector determined depth of shower maximum X using three different extensive air-shower (EAS) models – Sibyll2.3c [30], EPOS-LHC [31], and QGSJetII-04 [32]. These are referred to in the following figures, tables, and in combination with magnetic field models as SIBYLL, EPOS, and QGS4 for brevity.
@@ 3 Propagation Framework
The open-source CRPropa3 (v3.2) framework [26] is used to simulate UHECR primary nuclei propagation through intergalactic media. This software models interactions with cosmic backgrounds and deflections in magnetic fields, allowing us to trace the trajectories, composition, and energy changes of UHECRs from their sources to Earth. The five primary nuclei of hydrogen, helium, nitrogen, silicon, and iron are generated with an emission energy spectrum E, with energies ranging from 1018.6 eV to 1021 eV. This = 1 spectrum is generated for more statistics at the highest energies, and simulated events are reweighted for the combined fit to the data.
Simulated UHECR interactions with the intergalactic cosmic microwave, UV, optical, IR (Gilmore12 model [33]), and radio (Protheroe96 model [34]) photon backgrounds account for photopion production, nuclear photodisintegration, Bethe-Heitler process, gamma-gamma pair-production (single, double, and triplet), and inverse Compton scattering. Additionally, universe expansion adiabatic energy loss and unstable nuclear decays are considered.
The CRPropa3 observer used is a 200 kpc sphere centered at the Milky Way center and intersecting nuclei are recorded for energies E 1018.6 eV. Additional interactions and deflections in the last 200 kpc are expected to be negligible. Particles intersecting with another sphere 2 Gpc in radius or traveling further than 4 Gpc were discarded. Secondary neutrinos with energies above 100 TeV that intersect the observer are recorded. Secondary photons with energies above 100 MeV are immediately recorded if a projected linear path intersects the observer and are then propagated using the DintElecaPropagation module [35, 36].
@@ 3.1 Magnetic Fields
Charged nuclei deflections were simulated using four magnetic field models and a no magnetic field scenario. Two fields are turbulent and homogeneous with a mean strength B = 1 nG (BRMS = 1.1 nG) and length scales from 60 kpc to 1 Mpc (or 3 Mpc) and a Kolmogorov power spectrum for an average correlation length of = 234 kpc referred to as Rand.A (or = 647 kpc, Rand.B). The third model uses the Dolag et al. structured field [22], while the fourth employs the Hackstein et al. (CLUES) ’astrophysical1R’ structured field [23]. Figure 1 shows the magnetic field strength distributions. This sampled range of magnetic fields are within theoretical expectations and were chosen based on their prevalence in recent UHECR studies and their expected impact on the propagation outcomes.
The Dolag et al. field [22] was created via a magnetohydrodynamic simulation of the primordial evolution of the local universe in a cube 340 Mpc on a side (repeated in simulation as needed) and has structures that are well corresponded with nearby galaxy clusters (Coma, Virgo, Centaurus, Hydra, Perseus, and A3627). The field also includes large-scale structures such as sheets and filaments. Though the average magnetic field is small (B = 0.047 nG), the largest fields are on the order of 10 G. Though these high magnitude fields are a small proportion of the volume, this results in the large BRMS = 11 nG. The reported expectation is that ultra-high-energy protons will, on average, have a smaller deflection than experimental sensitivities [22].
@@ 4 Simulating FR0s as UHECR Sources
The simulated set of FR0s is upsampled from the well-sampled sky portion of the FR0CAT catalog [3] (limited to z 0.05) and extrapolated to a redshift z0.2 distance. This area is the spherical rectangle in supergalactic coordinates of -45∘ to 45∘ SGB (latitude) and 60∘ to 120∘ SGL (longitude) (for a solid angle ), as shown in Figure 2 on the left. Inside this 11.79% sky coverage there are 76 cataloged FR0s within z 0.05 that appear isotropically distributed. Upsampling these data results in 645 FR0s, also with isotropically distributed pointing-directions within the same redshift limit (Figure 2, right) and a total of 18,400 sources in an individual simulation within the extrapolated distance of z 0.2.
The simulated isotropic pointing-direction FR0s match the redshift distance distribution of data shown in Figure 3a, therefore the overall spatial distribution is not exactly volumetrically isotropic. The relative emission flux distribution is proportional to the radio output distribution (Figure 3b). Furthermore, the simulated FR0s preserve the correlation between radio (a source jet power measure)/UHECR flux and redshift (Figure 3c) to model the local universe source evolution.
As this method is probabilistic based on observational FR0 data, it is more realistic than the usual functional source evolution form used for combined fits of (1+z)m [37, 38, 39]. A fit to the emissivity versus redshift to this functional form results in . The cosmology used to convert FR0 catalog redshifts to distance was the Planck 2015 results [40].
The UHECR flux of each simulated FR0 source is emitted isotropically. The isotropic emission model for FR0s is based on their morphological characteristics approximating that they appear to possess relatively slow bulk flows, which suggest a lack of strong directional biases in their emissions. To refine our understanding of FR0 contributions, future work may explore anisotropic emission scenarios using emission cones and pointing-directions, assessing their impacts on UHECR propagation and detection.
@@ 4.1 Simulated Energy Spectrum
The simulated UHECR emission energies are a uniform distribution following an E energy spectrum for 1018.6 E/eV 1020.2. This spectrum is modified by interactions in the intergalactic medium during propagation to the observer, as demonstrated for the case with no magnetic field and an unmodified E emission spectrum in Figure 4. This propagated spectrum is modified when fitting to data by weighting the individual detected nuclei.

### v2

# Combined Fit of Spectrum and Composition for FR0 Radio-galaxy-emitted Ultra-high-energy Cosmic Rays with Resulting Secondary Photons and Neutrinos

@@ Abstract
This study comprehensively investigates the gamma-ray dim population of Fanaroff-Riley Type 0 (FR0) radio galaxies as potentially significant sources of ultra-high-energy cosmic rays (UHECRs, E 1018 eV) detected on Earth. While individual FR0 luminosities are relatively low compared to the more powerful Fanaroff-Riley Type 1 and Type 2 galaxies, FR0s are substantially more prevalent in the local universe, outnumbering the more energetic galaxies by a factor of 5 within a redshift of 0.05.
Employing CRPropa3 simulations, we estimate the mass-composition and energy spectra of UHECRs originating from FR0 galaxies for energies above 1018.6 eV. This estimation fits data from the Pierre Auger Observatory (Auger) using three extensive air shower models; both constant and energy-dependent observed elemental fractions are considered. The simulation integrates an approximately isotropic distribution of FR0 galaxies, extrapolated from observed characteristics, with UHECR propagation in the intergalactic medium, incorporating various plausible configurations of extragalactic magnetic fields, both random and structured. We then compare the resulting emission spectral indices, rigidity cutoffs, and elemental fractions with recent Auger results. In total, 25 combined energy-spectrum and mass-composition fits are considered.
Beyond the cosmic-ray fluxes emitted by FR0 galaxies, this study predicts the secondary photon and neutrino fluxes from UHECR interactions with intergalactic cosmic photon backgrounds. The multimessenger approach, encompassing observational data and theoretical models, helps elucidate the contribution of low-luminosity FR0 radio galaxies to the total cosmic-ray energy density.
@@
Extragalactic magnetic fields (507), Fanaroff-Riley radio galaxies (526), Gamma-rays (637), Low-luminosity AGN (2033), Neutrino astronomy (1100), Ultra-high-energy-cosmic radiation (1733)1 Introduction
@@
Ultra-high-energy cosmic rays (UHECRs) are the most energetic particles observed in the Universe, with energies exceeding eV. Their origins are an open astrophysical question, with potential sources including active galactic nuclei (AGN), various radio galaxies, gamma-ray bursts, and starburst galaxies. The Pierre Auger Observatory (Auger), located in Argentina, was established as the largest (3000 km2) UHECR hybrid detector in the world to study these particles [1]. Auger provides the most extensive statistics on the energy spectra, composition, and arrival directions of UHECR, offering essential information to explore their sources and propagation mechanisms.
@@
Low-luminosity Fanaroff-Riley (FR0-class radio-loud jetted AGN) radio galaxies [2] have emerged as a promising candidate class of sources of UHECRs detected on Earth. Recent surveys have highlighted that FR0s are approximately 5 times more prevalent in the local universe compared to their more energetic counterparts, Fanaroff-Riley Type 1 (FR1) and Type 2 (FR2) sources [3], qualifying them as potentially substantial contributors to the UHECR energy density. Despite their similar core radio properties to FR1s [4], FR0s display much weaker extended radio emission; however, their estimated jet power, which significantly exceeds the required cosmic-ray power per source [5], suggests they could substantially contribute to the UHECR flux measured on Earth. Studies utilizing the average FR0 spectral energy distribution suggest that these galaxies can accelerate UHECRs to the highest observed energies through hybrid acceleration, combining Fermi-I pre-acceleration with gradual shear acceleration [6]. Therefore, they fulfill both the Hillas criterion [7] and the energetics criteria [8, 9].
@@
Multimessenger astrophysics, which involves the combined study of various cosmic messengers such as photons, neutrinos, and cosmic rays, offers an even more comprehensive understanding of astrophysical phenomena. This approach helps identify and characterize sources of UHECRs and refine our knowledge of the mechanisms that generate and propagate these particles. Regarding the multimessenger properties of FR0s, lepto-hadronic jet-disc modeling of core-emission from quiet low-luminosity AGN jets predicts that FR0s are expected to be relatively weak gamma-ray and neutrino emitters [10, 11]. This theoretical evaluation of weak multimessenger emission aligns with the results of searches for gamma-ray-emitting FR0 galaxies. For example, stacking analyses were necessary to detect a very small unresolved FR0 population in Fermi-LAT data and identify them as gamma-ray emitters [12, 13]. A recent study [14] using various Fermi-LAT catalogs as sources of cosmic rays in CRPropa simulations concluded that previously resolved (therefore overwhelmingly non-FR0) gamma-ray sources cannot account for the large-scale anisotropy of UHECR events above 8 EeV as established by the Pierre Auger Collaboration [15]. This discrepancy led the authors of [14] to conclude that another population of UHECR sources, either lacking gamma-ray emission or unresolved by current-generation telescopes, must exist.
@@
The effect on expected arrival distribution of diffuse UHECRs in the intergalactic magnetic field has been examined in several studies, e.g., in [16, 17, 18] that provide constraints on the source emission spectra and composition (e.g., [19, 20]). The Auger large angular scale arrival direction analysis points to the extragalactic origins of UHECR for energies E 8 EeV [15]. A recent three-component combined-fit of the UHECR energy spectrum, extensive air shower (EAS) maxima distributions (X), and arrival directions supports an astrophysical model primarily comprised of homogeneous background sources and an adaptable addition from nearby source candidates [21]. For instance, if these additional nearby sources are a selection of 44 starburst galaxies, they could contribute up to a 20% flux fraction at 40 EeV [21]—a sizable fraction of the remaining bulk of UHECRs may be from FR0s.
@@
In this study, we leverage the strengths of multimessenger observations to investigate the properties of FR0 radio galaxies as potential UHECR sources. We simulate isotropic FR0 UHECR emission, extrapolating from measured FR0 properties and propagating them to Earth through plausible intergalactic magnetic fields using CRPropa3 [26]. By a combined fit of these simulations to energy and composition using Auger data [27, 28], we estimate the UHECR mass-composition and energy spectra emitted by FR0s. These combined energy-spectrum and composition fits are done using two nuclei emission models—a constant nuclei composition emission over the energy range and an evolving emission that depends on energy. Additionally, we present estimates of the cosmogenic secondary photon and neutrino fluxes resulting from UHECR interactions with cosmic photon backgrounds.
@@ 2 Pierre Auger Observatory Data
Auger performs high-precision and large-statistics analyses of energy spectra, mass-composition, and arrival directions of UHECR data which are essential for assessing theoretical models of cosmic-ray origins and propagation. This study utilizes the energy-spectrum and composition datasets to systematically extrapolate the observed characteristics of UHECRs to predictions of FR0 galaxy emission. The Auger data used in fitting simulations of FR0 radio galaxy UHECR emission is the publicly available energy spectrum from [28] and composition from [27] (using the method of [29]) for energies E 1018.6 eV. These results using 13 years of data are discussed further in Section 6.
The composition is given in terms of mean log mass number , which is evaluated from fluorescence-detector-determined depth of shower maximum X using three different extensive air-shower (EAS) models—Sibyll2.3c [30], EPOS-LHC [31], and QGSJetII-04 [32]. These are referred to in the following figures, tables, and in combination with magnetic field models as SIBYLL, EPOS, and QGS4 for brevity. UHECR data is an indirect experimental measurement of primary particle interaction, and their composition must be interpreted statistically through these EAS models that extrapolate the behavior of particle interactions in the atmosphere at much higher energies than available through direct experiments. The three most common EAS models are used in this study—the main two (EPOS and SIBYLL) are presented in figures for some clarity.
@@ 3 Propagation Framework
The open-source CRPropa3 (v3.2) framework [26] is used to simulate UHECR primary nuclei propagation through intergalactic media. This software models interactions with cosmic backgrounds and deflections in magnetic fields, allowing us to trace the trajectories, composition, and energy changes of UHECRs from their sources to Earth. The five primary nuclei of hydrogen, helium, nitrogen, silicon, and iron are generated with an emission energy spectrum E. This = 1 spectrum is generated for more statistics at the highest energies, and simulated events are reweighted for the combined fit to the data.
Simulated UHECR interactions with the intergalactic cosmic microwave, UV, optical, IR (Gilmore12 model [33]), and radio (Protheroe96 model [34]) photon backgrounds account for photopion production, nuclear photodisintegration, Bethe-Heitler process, gamma-gamma pair-production (single, double, and triplet), and inverse Compton scattering. Additionally, universe-expansion adiabatic energy loss and unstable nuclear decays are considered.
The CRPropa3 observer used is a 200 kpc sphere centered at the Milky Way center and intersecting nuclei are recorded for energies E 1018.6 eV. Additional interactions and deflections in the last 200 kpc are expected to be negligible. To conserve computation time, particles that intersected with another sphere of 2 Gpc in radius or traveled more than 4 Gpc were discarded, as such distances are unachievable within the universe’s lifetime. Secondary neutrinos with energies above 100 TeV that intersect the observer are recorded. Secondary photons with energies above 100 MeV are immediately recorded if a projected linear path intersects the observer and are then propagated using the DintElecaPropagation module [35, 36].
@@ 3.1 Magnetic Fields
Charged nuclei deflections were simulated using four magnetic field models and a no-magnetic-field scenario. Two fields are turbulent and homogeneous with a mean strength B = 1 nG (BRMS = 1.1 nG) and length scales from 60 kpc to 1 Mpc (or 3 Mpc) and a Kolmogorov power spectrum for an average correlation length of = 234 kpc referred to as Rand.A (or = 647 kpc, Rand.B). The third model uses the Dolag et al. structured field [22], while the fourth employs the Hackstein et al. (CLUES) ’astrophysical1R’ structured field [23]. Figure 1 shows the magnetic field strength distributions. This sampled range of magnetic fields are within theoretical expectations and were chosen based on their prevalence in recent UHECR studies and their expected impact on the propagation outcomes.
The Dolag et al. field [22] was created via a magnetohydrodynamic simulation of the primordial evolution of the local universe in a cube 340 Mpc on a side (repeated in simulation as needed) and has structures that are well corresponded with nearby galaxy clusters (Coma, Virgo, Centaurus, Hydra, Perseus, and A3627). The field also includes large-scale structures such as sheets and filaments. Though the average magnetic field is small (B = 0.047 nG), the largest fields are on the order of 10 G. Though these high-magnitude fields are a small proportion of the volume, this results in the large BRMS = 11 nG. The reported expectation is that ultra-high-energy protons will, on average, have a smaller deflection than experimental sensitivities [22].
@@ 4 Simulating FR0s as UHECR Sources
The simulated set of FR0s is upsampled from the well-sampled sky portion of the FR0CAT catalog [3] (limited to 0.05) and extrapolated to a redshift 0.2 distance. This area is the spherical rectangle in supergalactic coordinates of -45∘ to 45∘ SGB (latitude) and 60∘ to 120∘ SGL (longitude) (for a solid angle ), as shown in Figure 2 on the left. Inside this 11.79% sky coverage there are 76 cataloged FR0s within 0.05 that appear isotropically distributed. Upsampling these data results in 645 FR0s, also with isotropically distributed pointing-directions within the same redshift limit (Figure 2, right) and a total of 18,400 sources in an individual simulation within the extrapolated distance of 0.2.
The simulated isotropic-pointing-direction FR0s match the redshift distance distribution of data shown in Figure 3(a), therefore the overall spatial distribution is not exactly volumetrically isotropic. The relative emission flux distribution is proportional to the radio output distribution (Figure 3(b)). Furthermore, the simulated FR0s preserve the correlation between radio (a source jet power measure)/UHECR flux and redshift (Figure 3(c)) to model the local universe source evolution.
As this method is probabilistic based on observational FR0 data, it is more realistic than the usual functional source evolution form used for combined fits of (1+)m [37, 38, 39]. A fit to the emissivity versus redshift to this functional form results in . The cosmology used to convert FR0 catalog redshifts to distance was the Planck 2015 results [40].
The UHECR flux of each simulated FR0 source is emitted isotropically. The isotropic emission model for FR0s is based on their morphological characteristics approximating that they appear to possess relatively slow bulk flows, which suggest a lack of strong directional biases in their emissions. Simulating directed emission is not expected to have a significant effect due to the large number of sources and the necessary assumption of isotropically distributed emission-cone pointing-directions. To refine our understanding of FR0 contributions, future work may however explore anisotropic emission scenarios, assessing their impacts on UHECR propagation and detection.
@@ 4.1 Simulated Energy Spectrum
The simulated UHECR emission energies follow a uniform distribution with an E energy spectrum for energies 1018.6 E0/eV 1020.2. This spectrum is modified by interactions in the intergalactic medium during propagation to the observer, as demonstrated for the no-magnetic-field case with an unmodified E emission spectrum in Figure 4. This propagated spectrum is modified when fitting to data by weighting the individual detected nuclei.
The upper bound of E0 1020.2 eV was chosen because it is within the FR0 source-modeled maximum limit for iron nuclei given in [6], while the highest energy event observed by the Pierre Auger Observatory is currently 1020.2 eV [41]. It is assumed that events with such extreme observed energies may originate from an exceptional non-FR0 source class.
The best-case scenario in [6] (B-2) suggests a maximal mean rigidity of 1019.2 V. This implies a maximal energy of 1020.6 eV for iron, which is very close to 1020.2 eV. The maximal energy in model B-2 is determined by the size and magnetic field strength of the emission region. It is likely that other reasonable combinations of parameter values could lead to even higher energies. Crucially, the maximal energies provided in the models of [6] are exponential cutoff energies, meaning particles can be further accelerated. This cutoff energy is modeled by the rigidity-dependent exponential cutoff fit parameter discussed in Section 5.1.
As shown, the results from the combined-fit simulation can help inform future source-modeling efforts.